OpenStudy (anonymous):

One side of a triangle is three times the shortest side. The third side is 6 inches more than the shortest side. The perimeter of the triangle is 86 inches. Find the lengths of all three sides.

OpenStudy (anonymous):

@AlexandervonHumboldt2

Vocaloid (vocaloid):

first step is to define a variable let's let x = the shortest side with me so far?

OpenStudy (anonymous):

yes @Vocaloid

Vocaloid (vocaloid):

good. it also says that "one side is 3 times the shortest side" so we can write this as 3x with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

ok, it also says that The third side is 6 inches more than the shortest side. so third side = x + 6 with me so far?

OpenStudy (anonymous):

yes

OpenStudy (princeharryyy):

let shortest side = a other side = 3a third side = a +6 perimeter => a +3a + a +6 => 5a+6 = 86 5a = 80 a= 16(shortest) other 16*3 = 48 third one = 16+6 = 22

OpenStudy (princeharryyy):

but, I wonder if the triangle exists. hehehe :D :p

OpenStudy (anonymous):

is this right though? @princeharryyy

OpenStudy (princeharryyy):

yeah! it is. but, such a triangle is not possible as the axioms say.